# Marginal Metals and Kosterlitz-Thouless Type Phase Transition in Disordered Altermagnets

### v1

@@ Abstract
Altermagnetism, a recently discovered magnetic phase characterized by spin-split bands without net magnetization, has emerged as promising platform for novel physics and potential applications. However, its stability against disorder—ubiquitous in real materials—remains poorly understood. Here, we study the electron localization properties of two-dimensional altermagnets subject to disorder. Remarkably, we discover a disorder-driven phase transition from a marginal metallic phase to an insulator, which falls into the Kosterlitz-Thouless class. We demonstrate this by convincing numerical evidence and propose a vortex-antivortex-like spin pair picture for its interpretation. Moreover, we show that the characteristic spin anisotropy of altermagnets persists but gradually fades away across the transition. These changes directly affect the spin-splitting features that are detectable in angle-resolved photoemission spectroscopy and tunneling magnetoresistance. Our findings provide a new perspective on recent experimental observations of altermagnetism in candidate materials.
@@
Introduction.—Altermagnetism is an emerging collinear magnetic phase that uniquely combines advantages of ferromagnetism and antiferromagnetism Naka et al. (2019); Hayami et al. (2019); Šmejkal et al. (2022a, b); Bai et al. (2024); Song et al. (2025). In momentum space, altermagnets (AMs) exhibit highly anisotropic, non-relativistic spin splitting of electronic bands, e.g., manifesting $d$-wave symmetric patterns. From a real-space perspective, this distinctive feature arises because sublattices with opposite spins are related by rotational or mirror symmetry, rather than inversion or translation symmetry Šmejkal et al. (2022a, b). These combined properties renders AMs a fertile platform for exploring novel physics and potential applications in diverse fields such as spintronics, superconductivity, and topological phases González-Hernández, Rafael and Šmejkal, Libor and Výborný, Karel and Yahagi, Yuta and Sinova, Jairo and Jungwirth, Tomáš and Železný, Jakub (2021); Shao et al. (2021); Šmejkal et al. (2022c); Zhu et al. (2023); Zhang et al. (2024); Li and Liu (2023); Sun et al. (2023); Papaj (2023); Beenakker and Vakhtel (2023); Ouassou et al. (2023); Šmejkal et al. (2023); Liu et al. (2024a); Ghorashi et al. (2024); Zhou et al. (2024); Leeb et al. (2024); Sato et al. (2024); Sun et al. (2025); Duan et al. (2025); Gu et al. (2025); Lin et al. (2025); Antonenko et al. (2025); Hu et al. (2025a); Chen et al. (2025). Notably, altermagnetism has been reported in an increasing number of quantum materials Lee et al. (2024); Krempaský et al. (2024); Liu et al. (2024b); Zhu et al. (2025a); Zhang et al. (2025); Ahn et al. (2019); Yuan et al. (2020); Šmejkal et al. (2020); Ma et al. (2021a); Jiang et al. (2025); Reimers et al. (2024); Ding et al. (2024); Zeng et al. (2024); Yang et al. (2025); Liu et al. (2024c); Das et al. (2024); Xun et al. (2025).
@@
Despite rapidly growing interest in altermagnetism, the impact of disorder, which is inevitably present in real materials, on its stability is barely explored. To date, only a few works have studied the effects of single impurities in AMs Chen et al. (2024); Sukhachov and Linder (2024); Hu et al. (2025b); Maiani and Souto (2025); Gondolf et al. (2025). A crucial open question is thus: What is the fate of the AM phase in presence of disorder? This issue is particularly relevant in light of recent experiments. For example, in confirmed AM candidates at high doping levels, such as $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ Zhang et al. (2025), spin-1/2 dopant atoms may couple opposite electron spins, thereby potentially perturbing the underlying AM order. Moreover, many proposed AM materials, including RuO2 Ahn et al. (2019); Yuan et al. (2020); Šmejkal et al. (2020), $\mathrm{KV_{2}Se_{2}O}$ Ma et al. (2021a); Jiang et al. (2025) and $\mathrm{CrSb}$ Reimers et al. (2024); Ding et al. (2024); Zeng et al. (2024); Yang et al. (2025), are metallic. In such metallic systems, disorder could significantly alter electronic states and transport properties.
@@
Figure 1: (a) Momentum-resolved local density of states (LDOS) of the $d$-wave AM at disorder strength $W=2t$. (b) Normalized localization length $\lambda/L$ as a function of $W$ for increasing system width $L$. $\lambda$ is calculated on a long ribbon with width $L$ and length of $3\times 10^{6}$. (c) Single-parameter fitting of the correlation length $\xi$ near the KT-type phase transition, extracted from the data in (b). On the insulating side, $\xi$ scales as $\ln\xi\propto|W-W_{c}|^{-1/2}$, where $W_{c}$ denotes the critical disorder strength. Inset: Collapse of the data from (b) into a single curve under finite-size scaling. (d) Phase diagram in the $W$-$t_{J}$ plane. Other parameters are $t_{J}=0.3t$ and $\mu=-2t$.
@@
In this work, we systematically investigate how disorder affects the stability of the AM phase in two dimensions (2D). Strikingly, we find that metallic AMs persist as an exotic marginal metallic state with zero net magnetization over a finite range of disorder strengths, even when scattering between opposite spins becomes significant. Upon further increasing disorder strength, the AMs undergo a distinct metal–insulator phase transition. Through extensive scaling analyses of correlation length and conductance in a representative $d$-wave AM system, we find that this phase transition belongs to the Kosterlitz-Thouless (KT) class. To understand this phase transition, we provide a physical picture based on vortex-antivortex-like spin pairs. Furthermore, we find that the characteristic spin anisotropy of the AM persists but gradually diminishes during the transition. This behavior can be probed experimentally via angle-resolved photoemission spectroscopy (ARPES) and tunneling magnetoresistance. Our results unveil the particular impact of disorder on AMs and suggest disorder as a potential reason for the difficulty in detecting the predicted spin splitting in candidate materials such as RuO2.
@@
Model and general analysis.Altermagnets feature alternating spin splitting in electronic bands, dictated by spin-space group symmetries Šmejkal et al. (2022b); Liu et al. (2022). To elucidate the essential physics, we consider a disordered $d$-wave AM described by the minimal tight-binding Hamiltonian on a square lattice Šmejkal et al. (2022a, b)
@@
When disorder is absent, the two spins are decoupled. The system respects $\text{[}C_{2}||C_{4z}]$ spin symmetry, i.e., a fourfold spatial rotation about $z$-axis combined with a spin flip. This symmetry imposes $d$-wave altermagnetism with zero net magnetization. The two spin-polarized bands are given by $E_{\uparrow(\downarrow)}({\bf k})=-2t(\cos k_{x}+\cos k_{y})\mp 2t_{J}(\cos k_{x}-\cos k_{y})-\mu$, where ${\bf k}=(k_{x},k_{y})$ is the 2D momentum and the lattice constant is set to $a=1$. It gives rise to anisotropically spin-split Fermi surfaces, as illustrated in Fig. 1(a).
@@
We consider onsite disorder [the last term in Eq. (1)], which can couple the spin degrees of freedom. Specifically, the $w_{i}^{0}\sigma_{0}$ term represents nonmagnetic disorder, while $w_{i}^{x}\sigma_{x}$ and $w_{i}^{y}\sigma_{y}$ correspond to magnetic disorder that flip the spins Šmejkal et al. (2022c); Li2 (a). Here, $w_{i}^{s}$ are random values taken from a uniformly distributed range $[-W/2,W/2]$, with $W$ the disorder strength. Note that this model explicitly breaks time reversal symmetry, placing it in the unitary class (class $A$) in 2D Altland and Zirnbauer (1997); Schnyder et al. (2008). Generally, electronic states in this class will be localized even at infinitesimal disorder strength, except for quantum Hall plateau transitions Wei et al. (1988); Ando (1983). Surprisingly, as we show below, the AM metals exhibit particular extended states even under strong disorder, defying this conventional expectation.
@@
KT-type transition in disordered AM.—To study the stability of AM metals under disorder, we consider a long ribbon geometry with width $L_{y}=L$ and length $L_{x}\gg L$. Whether the states remain delocalized or become localized can be deduced from the localization length under finite-size scaling. We compute the localization length $\lambda$ by means of transfer matrix MacKinnon and Kramer (1983); Kramer and MacKinnon (1993); Yamakage et al. (2013), and define a normalized localization length as $\Lambda\equiv\lambda(L,W)/L$. According to finite-size scaling theory, $\Lambda$ increases with $L$ for metallic states, decreases with $L$ for insulating states, and remains scale-invariant at a critical point MacKinnon and Kramer (1983).
@@
Figure 1(b) shows the scaling behavior of $\Lambda$ in the disordered AM. Remarkably, over a wide range of disorder strengths, $\Lambda$ remains independent of $L$, indicating a marginal metallic state. We refer to this phase as the altermagnetic marginal metal (AMMM). Its emergence suggests that the AM metallic state is resilient to disorder up to a considerable strength, which is crucial for the transport characteristics discussed later. Beyond a critical disorder strength, however, $\Lambda$ decreases monotonically as $L$ grows, indicating that the system is turned into a localized state. It thus realizes a disorder-driven metal-insulator transition. The appearance of the scale-invariant AMMM phase implies that this transition belongs to KT class.
@@
with fitting parameters $b=9.24\pm 0.21$ and $W_{c}=4.13t\pm 0.08t$. Moreover, all finite-size scaling curves for different $L$ collapse to a single curve [inset of Fig. 1(c)]. These results strongly support that the metal-insulator transition is of KT type  [^1] For KT transition in 2D XY model, the correlation length scales as $\xi(T)\propto\exp[b/\sqrt{T-T_{c}}]$ above $T_{C}$ with $T$ the temperature.. Additionally, by repeating the above analysis for varying values of $t_{J}$, we map out the phase diagram in the $t_{J}$-$W$ plane, shown in Fig. 1(d). The diagram shows that $W_{c}$ increases with $t_{J}$, indicating enhanced robustness of the metallic phase by stronger altermagnetism.
@@
Spectral statistics.—For a comprehensive understanding of disordered AMs, we examine the spectral statistics using level spacing ratio (LSR). The LSR is defined as $r_{n}=\frac{\mathrm{min}(s_{n},s_{n-1})}{\mathrm{max}(s_{n},s_{n-1})}$, where $s_{n}\equiv E_{n+1}-E_{n}$ is the spacing between adjacent energy levels $E_{n}$ and $E_{n+1}$ Oganesyan and Huse (2007). Figure S3(a) plots disorder-averaged LSR $\langle r\rangle$ against disorder strength $W$. For a finite range of $W<W_{c}$, $\langle r\rangle$ remains nearly constant at $\langle r\rangle\approx 0.6$, consistent with the metallic nature of AMMMs ($\langle r\rangle_{M}=0.6$). As $W$ exceeds $W_{c}$, $\langle r\rangle$ decreases smoothly toward $\langle r\rangle_{I}=0.386$, the universal value characteristic of localized insulating states Atas et al. (2013). The probability distributions $P(r)$ at representative $W$ are shown in Fig. S3(b). For weak disorder ($W=3t$), $P(r)$ closely follows the Gaussian unitary ensemble (GUE) form $P_{\text{GUE}}(r)=\frac{81\sqrt{3}}{2\pi}\frac{(r+r^{2})^{2}}{(1+r+r^{2})^{4}}$ Atas et al. (2013), which is typical of extended states. For strong disorder ($W=14t$), $P(r)$ instead matches the Poisson distribution $P_{p}(r)=\frac{2}{(1+r)^{2}}$, indicative of uncorrelated, localized states. In the SM, we also analyze inverse participation ratio statistics Li2 (b). These results further confirm the existence of a robust AMMM phase and the disorder-induced metal–insulator transition.
@@
Figure 3: (a) Averaged LSR $\langle r\rangle$ as a function of $W$. (b) Distribution of LSR in the metallic and insulating limits.Other parameters are as the same as Fig. 1.
@@
Phenomenological picture.—We present a phenomenological argument to interpret the KT-type phase transition in disordered AMs. In the 2D XY model, the KT transition arises from the unbinding of vortex-antivortex pairs as temperature increases Kosterlitz and Thouless (1973). Analogously, the spin-up and spin-down degrees of freedom in disordered AMs can be viewed as vortex-like and antivortex-like excitations, respectively. Weak magnetic disorder, described by the terms $w_{i}^{x}\sigma_{x}$ and $w_{i}^{y}\sigma_{y}$, couples spin-up and spin-down states (quantized along $\sigma_{z}$), effectively forming spin-paired states ($\bm{\uparrow}\leftrightsquigarrow\bm{\downarrow}$) akin to bound vortex-antivortex pairs thermally activated in the XY model. Crucially, the $[C_{2}||C_{4z}]$ spin symmetry (equivalently, the $C_{4}\mathcal{T}$ symmetry that combines four-fold rotation and time-reversal symmetries) protects the marginal metallic phase Wang et al. (2024). Indeed, as shown in Fig. 1(d), the critical disorder strength $W_{c}$ increases with the AM strength $t_{J}$. In absence of AM ($t_{J}=0$) or in conventional antiferromagnets, no marginal metal phase emerges, underscoring the essential role of altermagnetism. When the disorder strength exceeds $W_{c}$, magnetic scattering between opposite spins becomes so strong that the spin pairs are fully depaired, resembling the breakdown of vortex-antivortex pairs in the XY model. The system then enters an insulating state. In this way, the AM undergoes a KT-type phase transition driven by disorder.
@@
Figure 4: (a) Spin-resolved LDOS $A_{\sigma}({\bf k})$ (left and middle) and total LDOS $A({\bf k})=A_{\uparrow}({\bf k})+A_{\downarrow}({\bf k})$ (right) in momentum space for disorder strength $W=t$. (b) The same as (a) but for $W=6t$. Other parameters are $t_{J}=0.3t,$ $\mu=-2t$, and $\eta=10^{-3}t$.
@@
Mitigation of spin anisotropy by disorder.—Disorder significantly affects the characteristic spin anisotropy of AMs, namely, the alternating spin splitting in momentum space. To illustrate this, we track the evolution of spin anisotropy during the phase transition by analyzing spin-resolved LDOS. The LDOS is calculated from the disorder-averaged effective Hamiltonian $\langle H({\bf k})\rangle$ via
@@
where $\eta$ is a small spectral broadening Li2 (b). As shown in Fig. 4(a), the spin anisotropy remains prominent in the AMMM phase at weak disorder: spin-up and spin-down Fermi surfaces exhibit distinct orientations, confirming the persistence of spin-splitting. As disorder strength $W$ increases, the bands broaden due to disorder scattering, and the Fermi-surface splitting gets increasingly blurred [see Figs. 4(a,b) for comparison]. In this regime, the anisotropic $d$-wave AM gradually evolves towards an isotropic pattern. This transition process is similar to the disorder-driven suppression of $d$-wave superconducting pairing Balatsky et al. (2006); Spivak et al. (2009). When $W$ exceeds $W_{c}$, spin anisotropy rapidly fades away [Fig. 4(b)]. In this strong-disorder regime, spin-up and spin-down states strongly scatter into each other, and the system eventually enters an insulating phase. Note that throughout this evolution, the system maintains zero net magnetization, and spin remains a good quantum number after disorder average.
@@
Figure 5: (a) Schematic of the tunneling junction composed of two AM layers of length $L_{x}$ separated by an insulating barrier. The Néel vectors of two AM layers are configured to be either parallel (P) or antiparallel (AP). The system size is $L_{x}=10a$ and $L_{y}=40a$. (b) Tunneling conductances $G_{P}$ and $G_{AP}$ as functions of energy $E$ for the parallel and antiparallel configurations at $W=0$. (c) $G_{P}$ and $G_{AP}$ as functions of $W$ at different $E$. (d) TMR as a function of $W$. Inset: Energy spectrum of an AM ribbon of width $L_{y}=40$ along $x$ direction. Other parameters are as the same as Fig. 1.
@@
Tunneling magnetoresistance in disordered AMs.The spin splitting affected by disorder in AMs can also be detected via tunneling magnetoresistance in tunnel junctions. We consider a junction composed of two AM layers with either parallel or antiparallel Néel vectors, separated by a tunnel barrier, as illustrated in Fig. 5(a). Recent studies have shown that tunneling magnetoresistance naturally arises in such junctions owing to the intrinsic spin splitting Shao et al. (2021); Šmejkal et al. (2022c); Liu et al. (2024e); Noh et al. (2025). Within the energy range of the band width, we calculate the tunneling conductance using $G=\frac{e^{2}}{h}\mathrm{Tr}[TT^{\dagger}]$, where $T$ is the transmission matrix across the junction. Due to the spin anisotropy, the conductance for the parallel configuration $G_{P}$ is generally larger than that of the antiparallel configuration $G_{AP}$, i.e., $G_{P}>G_{AP}$. This feature is confirmed in Fig. 5(b), which shows $G_{P}$ and $G_{AP}$ as functions of energy. The peaks in $G_{P}(E)$ originate from tunneling resonances in the junction.

### v2

@@ Abstract
Altermagnetism, a recently discovered magnetic phase characterized by spin-split bands without net magnetization, has emerged as promising platform for novel physics and potential applications. However, its stability against disorder—ubiquitous in real materials—remains poorly understood. Here, we study the electron localization properties of two-dimensional $d$-wave altermagnets subject to disorder. Remarkably, we discover a disorder-driven phase transition from a marginal metallic phase to an insulator, which falls into the Kosterlitz-Thouless class. We demonstrate this by strong numerical evidence and propose an interpretation in terms of vortex-antivortex pairs in the disorder-induced local in-plane spin magnetization. Moreover, we show that the characteristic spin anisotropy of altermagnets persists but gradually fades away across the transition. These changes directly affect the spin-splitting features that are detectable in angle-resolved photoemission spectroscopy and tunneling magnetoconductance. Our findings provide a new perspective on recent experimental observations of altermagnetism in candidate materials.
@@
Introduction.—Altermagnetism is an emerging collinear magnetic phase that uniquely combines advantages of ferromagnetism and antiferromagnetism Šmejkal et al. (2020); Naka et al. (2019); Ahn et al. (2019); Hayami et al. (2019); Yuan et al. (2020); Šmejkal et al. (2022a, b); Bai et al. (2024); Song et al. (2025). In momentum space, altermagnets (AMs) exhibit highly anisotropic, non-relativistic spin splitting of electronic bands, e.g., manifesting $d$-wave symmetric patterns. From a real-space perspective, this distinctive feature arises because sublattices with opposite spins are related by rotational or mirror symmetry, rather than inversion or translation symmetry Šmejkal et al. (2022a, b). These combined properties renders AMs a fertile platform for exploring novel physics and potential applications in diverse fields such as spintronics, superconductivity, and topological phases R. González-Hernández, L. Šmejkal, Karel Výborný, Y. Yahagi, J. Sinova, T. Jungwirth, and J. Železný (2021); Shao et al. (2021); Šmejkal et al. (2022c); Zhu et al. (2023); Zhang et al. (2024); Li and Liu (2023); Sun et al. (2023); Papaj (2023); Beenakker and Vakhtel (2023); Ouassou et al. (2023); Šmejkal et al. (2023); Liu et al. (2024a); Ghorashi et al. (2024); Zhou et al. (2024); Leeb et al. (2024); Sato et al. (2024); Sun et al. (2025); Duan et al. (2025); Gu et al. (2025); Lin et al. (2025); Antonenko et al. (2025); Hu et al. (2025a); Chen et al. (2025). Notably, altermagnetism has been reported in an increasing number of quantum materials Lee et al. (2024); Krempaský et al. (2024); Liu et al. (2024b); Zhu et al. (2025a); Zhang et al. (2025a); Ahn et al. (2019); Yuan et al. (2020); Šmejkal et al. (2020); Ma et al. (2021a); Jiang et al. (2025); Zhang et al. (2025b); Reimers et al. (2024); Ding et al. (2024); Zeng et al. (2024); Yang et al. (2025); Liu et al. (2024c); Das et al. (2024); Xun et al. (2025).
@@
Despite rapidly growing interest in altermagnetism, the impact of disorder, which is inevitably present in real materials, on its stability is barely explored. To date, only a few works have studied the effects of single impurities in AMs Chen et al. (2024); Sukhachov and Linder (2024); Hu et al. (2025b); Maiani and Souto (2025); Gondolf et al. (2025). A crucial open question is thus: What is the fate of the AM phase in presence of disorder? This issue is particularly relevant in light of recent experiments. For example, in confirmed AM candidates at high doping levels, such as $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ Zhang et al. (2025a), spin-1/2 dopant atoms may couple opposite electron spins, thereby potentially perturbing the underlying AM order. Moreover, many proposed AM materials, including RuO2 Šmejkal et al. (2020); Ahn et al. (2019), $\mathrm{KV_{2}Se_{2}O}$ Ma et al. (2021a); Jiang et al. (2025); Zhang et al. (2025a) and $\mathrm{CrSb}$ Reimers et al. (2024); Ding et al. (2024); Zeng et al. (2024); Yang et al. (2025), are metallic. In such metallic systems, disorder could significantly alter electronic states and transport properties.
@@
Figure 1: (a) Momentum-resolved local density of states (LDOS) of the $d$-wave AM at disorder strength $W=2t$. (b) Normalized localization length $\lambda/L$ as a function of $W$ for increasing system width $L$. $\lambda$ is calculated on a long ribbon with width $L$ and length of $3\times 10^{6}a$. (c) Single-parameter fitting of the correlation length $\xi$ near the KT-type phase transition, extracted from the data in (b). On the insulating side, $\xi$ scales as $\ln\xi\propto|W-W_{c}|^{-1/2}$, where $W_{c}$ denotes the critical disorder strength. Inset: Collapse of the data from (b) into a single curve under finite-size scaling. (d) Phase diagram in the $W$-$t_{J}$ plane. Other parameters are $t_{J}=0.3t$ and $\mu=-2t$.
@@
In this work, we systematically investigate how disorder affects the stability of the AM phase in two dimensions (2D). Strikingly, we find that metallic AMs persist as an exotic marginal metal with zero net magnetization over a finite range of disorder strengths, even when opposite-spin scattering becomes significant. Beyond a critical disorder strength, the system undergoes a distinct metal–insulator transition. Through extensive scaling analyses of correlation length and conductance in a representative $d$-wave AM system, we find that this phase transition belongs to the Kosterlitz-Thouless (KT) class. To understand this transition, we propose a physical picture based on vortex-antivortex pairs in the induced local in-plane spin magnetization. Furthermore, the characteristic spin anisotropy persists at weak disorder but gradually diminishes across the transition. This behavior can be probed via angle-resolved photoemission spectroscopy (ARPES) and tunneling magnetoconductance. The results hold for both uncorrelated and correlated disorder and generalize to $g$-wave AMs and other lattices, for instance, the Lieb lattice. Our findings underscore the particular impact of disorder on AMs and suggest disorder as a potential reason for the difficulty in detecting the predicted spin splitting in candidate materials such as RuO2.
@@
Model and general analysis.Altermagnets feature alternating spin splitting in electronic bands, related to spin-space group symmetries Šmejkal et al. (2022b); Liu et al. (2022). For concreteness, we consider a disordered $d$-wave AM described by the minimal tight-binding Hamiltonian on a square lattice Šmejkal et al. (2022a, b)
@@
When disorder is absent, the two spins are decoupled. The system respects $\text{[}C_{2}||C_{4z}]$ spin symmetry, i.e., a fourfold spatial rotation about the $z$-axis combined with a spin flip. This symmetry imposes $d$-wave altermagnets described by $H_{0}({\bf k})=-(2t\cos k_{x}+2t\cos k_{y}+\mu)\sigma_{0}-2t_{J}(\cos k_{x}-\cos k_{y})\sigma_{z}$ with zero net magnetization. The two spin-polarized bands are given by $E_{\uparrow(\downarrow)}({\bf k})=-2t(\cos k_{x}+\cos k_{y})\mp 2t_{J}(\cos k_{x}-\cos k_{y})-\mu$, where ${\bf k}=(k_{x},k_{y})$ is the 2D momentum and the lattice constant is set to $a=1$. It gives rise to anisotropically spin-split Fermi surfaces, as illustrated in Fig. 1(a).
@@
We consider onsite disorder [the last term in Eq. (1)], which can couple the spin degrees of freedom. Specifically, the $w_{{\bf r}}^{0}\sigma_{0}$ term represents nonmagnetic disorder, while $w_{{\bf r}}^{x}\sigma_{x}$ and $w_{{\bf r}}^{y}\sigma_{y}$ correspond to magnetic disorder that flip the spins Šmejkal et al. (2022c) [^1] Note that the main results we discuss will remain if the disorder of form $w_{{\bf r}}^{z}\sigma_{z}$ is added. This type of disorder does not flip the spin. Here, $w_{{\bf r}}^{s}$ are random values taken from a uniformly distributed range [W/2,W/2][-W/2,W/2], with $W$ the disorder strength. Note that this model explicitly breaks time reversal symmetry, placing it in the unitary class (class $A$) in 2D Altland and Zirnbauer (1997); Schnyder et al. (2008). Generally, electronic states in this class will be localized even at infinitesimal disorder strength, except for quantum Hall plateau transitions Wei et al. (1988); Ando (1983). Surprisingly, as we show below, the AM metals exhibit particular extended states even under strong disorder, defying this conventional expectation.
@@
KT-type transition in disordered AM.—To study the stability of AM metals in presence of disorder, we consider a long ribbon geometry with width $L_{y}=L$ and length $L_{x}\gg L$. Whether the states remain delocalized or become localized can be deduced from the localization length under finite-size scaling. We compute the localization length $\lambda$ by means of the transfer matrix MacKinnon and Kramer (1983); Kramer and MacKinnon (1993); Yamakage et al. (2013), and define a normalized localization length as $\Lambda\equiv\lambda(L,W)/L$. According to finite-size scaling theory, $\Lambda$ increases with $L$ for metallic states, decreases with $L$ for insulating states, and remains scale-invariant at a critical point MacKinnon and Kramer (1983).
@@
Figure 1(b) shows the scaling behavior of $\Lambda$ in the disordered AM. Remarkably, over a wide range of disorder strengths, $\Lambda$ remains independent of $L$, indicating a marginal metallic state. We refer to this phase as the altermagnetic marginal metal (AMMM). Its emergence suggests that the AM metallic state is resilient to disorder up to a considerable strength, which is crucial for the transport characteristics discussed later. Beyond a critical disorder strength, however, $\Lambda$ decreases monotonically as $L$ grows, indicating that the system is turned into an insulator. It thus realizes a disorder-driven metal-insulator transition. The appearance of the scale-invariant AMMM phase implies that this transition belongs to the KT class.
@@
with fitting parameters $b=9.24\pm 0.21$ and $W_{c}=4.13t\pm 0.08t$. Moreover, all finite-size scaling curves for different $L$ collapse to a single curve [inset of Fig. 1(c)]. These results strongly support that the metal-insulator transition is of KT type  [^2] For KT transition in 2D XY model, the correlation length scales as $\xi(T)\propto\exp[b/\sqrt{T-T_{c}}]$ above $T_{C}$ with $T$ the temperature.. Additionally, by performing the above analysis for different values of $t_{J}$, we map out the phase diagram in the $t_{J}$-$W$ plane, shown in Fig. 1(d). The diagram shows that $W_{c}$ increases with $t_{J}$, indicating enhanced robustness of the metallic phase by stronger altermagnetism.
@@
Spectral statistics.—For a comprehensive understanding of disordered AMs, we examine the spectral statistics based on level spacing ratio (LSR). The LSR is defined as $r_{n}=\frac{\mathrm{min}(s_{n},s_{n-1})}{\mathrm{max}(s_{n},s_{n-1})}$, where $s_{n}\equiv E_{n+1}-E_{n}$ is the spacing between adjacent energy levels $E_{n}$ and $E_{n+1}$ Oganesyan and Huse (2007). Figure S4(a) plots the disorder-averaged LSR $\langle r\rangle$ against disorder strength $W$. For a finite range of $W<W_{c}$, $\langle r\rangle$ remains nearly constant at $\langle r\rangle\approx 0.6$, consistent with the metallic nature of AMMMs ($\langle r\rangle_{M}=0.6$). As $W$ exceeds $W_{c}$, $\langle r\rangle$ decreases smoothly toward $\langle r\rangle_{I}=0.386$, the universal value characteristic of localized insulating states Atas et al. (2013). The probability distributions $P(r)$ at representative $W$ are shown in Fig. S4(b). For weak disorder ($W=3t$), $P(r)$ closely follows the Gaussian unitary ensemble (GUE) form $P_{\text{GUE}}(r)=\frac{81\sqrt{3}}{2\pi}\frac{(r+r^{2})^{2}}{(1+r+r^{2})^{4}}$ Atas et al. (2013), which is typical of extended states. For strong disorder ($W=14t$), $P(r)$ instead matches the Poisson distribution $P_{p}(r)=\frac{2}{(1+r)^{2}}$, indicative of uncorrelated, localized states. In the SM Li (2), we also analyze inverse participation ratio statistics. These results further confirm the existence of a robust AMMM phase and the disorder-induced metal–insulator transition  [^3] Note that there exists a mobility edge in the whole spectrum, but we focus on the properties at the Fermi energy here.
@@
Figure 3: (a) Averaged LSR $\langle r\rangle$ as a function of $W$. (b) Distribution of LSR in the metallic and insulating limits. We take $2\times 10^{3}$ disorder configurations. Other parameters are as the same as Fig. 1.
@@
Phenomenological picture.—We present a phenomenological argument to interpret the KT-type phase transition in disordered AMs. In the 2D XY model, the KT transition arises from the unbinding of vortex-antivortex pairs as temperature increases Kosterlitz and Thouless (1973). Analogously, in 2D AMs, magnetic disorder induces spin flips that generate pronounced in-plane spin components, whose spatial texture admits vortices and antivortices Li (2). The disorder strength $W$ plays the role of temperature in the XY model. At weak $W$, we can directly visualize the appearance of in-plane spin vortex-antivortex pairs Li (2) corresponding to the marginal metallic phase. Note that the $C_{2}||C_{4z}]$ symmetry (equivalently, the $C_{4z}\mathcal{T}$ symmetry that combines four-fold rotation and time-reversal symmetries) can protect the marginal metallic phase Wang et al. (2024) [^4] The marginal metal phase can be removed by explicitly breaking $C_{4z}\mathcal{T}$, for instance, by adding a term $t_{e}(\cos k_{x}+\cos k_{y})\sigma_{z}$ to the Hamiltonian $H_{0}({\bf k})$. Indeed, as shown in Fig. 1(d), the critical disorder strength $W_{c}$ increases with the AM strength $t_{J}$. In absence of AM ($t_{J}=0$) or in conventional antiferromagnets, no marginal metal phase emerges, underscoring the essential role of altermagnetism. The marginal metallic phase corresponds to the quasi-ordered phase in the XY model at low temperatures. When the disorder strength exceeds $W_{c}$, magnetic scattering between opposite spins becomes so strong that the in-plane spin vortices proliferate and unpair Li (2), resembling the breakdown of vortex-antivortex pairs in the XY model at high temperatures. The system then enters an insulating state. In this way, the AM undergoes a KT-type phase transition driven by disorder.
@@
Figure 4: (a) Spin-resolved LDOS $A_{\sigma}({\bf k})$ (left and middle) and total LDOS $A({\bf k})=A_{\uparrow}({\bf k})+A_{\downarrow}({\bf k})$ (right) in momentum space for disorder strength $W=t$. (b) The same as (a) but for $W=6t$. We take $5\times 10^{4}$ disorder configurations. Other parameters are $t_{J}=0.3t,$ $\mu=-2t$, and $\eta=10^{-3}t$.
@@
Mitigation of spin anisotropy by disorder.Disorder significantly affects the characteristic spin anisotropy of AMs, namely, the alternating spin splitting in momentum space. To illustrate this, we track the evolution of spin anisotropy during the phase transition by analyzing spin-resolved LDOS. The LDOS is calculated from the disorder-averaged effective Hamiltonian $\langle H({\bf k})\rangle$ via
@@
where $\eta$ is a small spectral broadening Li (2). As shown in Fig. 4(a), the spin anisotropy remains prominent in the AMMM phase at weak disorder: spin-up and spin-down Fermi surfaces exhibit distinct orientations, confirming the persistence of spin-splitting. As the disorder strength $W$ increases, the bands broaden due to disorder scattering, and the Fermi-surface splitting gets blurred [see Figs. 4(a,b) for comparison]. In this regime, the anisotropic $d$-wave AM gradually evolves towards an isotropic pattern. This transition process is similar to the disorder-driven suppression of $d$-wave superconducting pairing Balatsky et al. (2006); Spivak et al. (2009). When $W$ exceeds $W_{c}$, spin anisotropy rapidly fades away [Fig. 4(b)]. In this strong-disorder regime, spin-up and spin-down states strongly scatter into each other, and the system eventually enters an insulating phase. Note that throughout this evolution, the system maintains zero net magnetization, and spin remains a good quantum number after disorder average.
@@
Figure 5: (a) Schematic of the tunneling junction composed of two AM layers of length $L_{x}$ separated by an insulating barrier. The Néel vectors of two AM layers are configured to be either parallel (P) or antiparallel (AP). The system size is $L_{x}=10a$ and $L_{y}=40a$. (b) Tunneling conductances $G_{P}$ and $G_{AP}$ as functions of energy $E$ for the parallel and antiparallel configurations at $W=0$. (c) $G_{P}$ and $G_{AP}$ as functions of $W$ at different $E$. (d) TMC as a function of $W$. Inset: Energy spectrum of an AM ribbon of width $L_{y}=40a$ along $x$ direction. We take $2\times 10^{4}$ disorder configurations. Other parameters are as the same as Fig. 1.
@@
Tunneling magnetoconductance in disordered AMs.The spin splitting affected by disorder in AMs can also be detected via tunneling magnetoconductance in tunnel junctions. We consider a junction composed of two AM layers with either parallel or antiparallel Néel vectors, separated by a tunnel barrier, as illustrated in Fig. 5(a). Recent studies have shown that tunneling magnetoresistance naturally arises in such junctions owing to the intrinsic spin splitting Shao et al. (2021); Šmejkal et al. (2022c); Liu et al. (2024e); Noh et al. (2025). Within the energy range of the band width, we calculate the tunneling conductance using $G=\frac{e^{2}}{h}\mathrm{Tr}[TT^{\dagger}]$, where $T$ is the transmission matrix across the junction. Due to the spin anisotropy, the conductance for the parallel configuration $G_{P}$ is generally larger than that of the antiparallel configuration $G_{AP}$, i.e., $G_{P}>G_{AP}$. This feature is confirmed in Fig. 5(b), which shows $G_{P}$ and $G_{AP}$ as functions of energy. The peaks in $G_{P}(E)$ originate from tunneling resonances in the junction.